Please show every step of the solutions. Thanks in advance!
Answer #1

Let the tension in the upper string be x and that in the lower string be y.

Balancing vertical forces,

x*1.5/2-y*1.5/2=mg

Balancing horizontal forces,

x*sqrt(2^2-1.5^2)/2+y*sqrt(2^2-1.5^2)/2=mv^2/sqrt(2^2-1.5^2)

Solving for x and y,

x=0.67mg+0.5714mv^2

y=0.5714mv^2-0.67mg

2)Assuming the car begins at rest,

speed at the end of the circle=sqrt(2*at*2pi*r)=2*sqrt(at*pi*r)

Radial acceleration=v^2/r=(2*at*2pi)=4(at*pi)

Angle with radius=Tan^-1(at*/4(at*pi))=Tan^-1(1/4(pi))=4.55 degrees